OpenStudy (anonymous):

What is the slope of this equation? y = 8x - 7

OpenStudy (anonymous):

if u could explain that would be helpfull

OpenStudy (anonymous):

This is in the form y = mx +b where m is the slope. What do you think is the slope?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

yup i am so stupid tahnk u

OpenStudy (anonymous):

No you are not. You figured it out. Good job.

OpenStudy (anonymous):

hey another question

OpenStudy (anonymous):

y=x-1

OpenStudy (anonymous):

i mean y=7

OpenStudy (anonymous):

You need the slope for that one too?

OpenStudy (anonymous):

in here its asking for the slope so ya

OpenStudy (anonymous):

y = 7 ??? it is just stating y = 7

OpenStudy (anonymous):

but y isnt the slope m is or k is

OpenStudy (anonymous):

If the question is y = 7 it is just stating y equals 7 there is no slope

OpenStudy (anonymous):

k

OpenStudy (anonymous):

if it is y = x -1 then the slope is 1 because x = 1x

OpenStudy (anonymous):

i am doing a online homework and they did not help me understand it so ya very appreciate u helping me

OpenStudy (anonymous):

No problem and you are welcome.

OpenStudy (anonymous):

Y = 2x + 4 is it a non-proportional or proportional linear relationship? how do i solve does

OpenStudy (anonymous):

*those